How many gallons of a 10% salt solution should be added to 20 gallons of a 60% salt solution to obtain a 50% salt solution?
Found 2 solutions by checkley77, dabanfield:
Answer by checkley77(12844)

.60*20+.10x=.50(20+x)
12+.10x=10+.50x
.10x-.50x=10-12
-.4x=-2
x=-2/-.40
x=5 gallons of 10% salt solution is added.
Proof:
.60*20+.10*5=.5(20+5)
12+.5=.5*25
12.5=12.5

Answer by dabanfield(803)

How many gallons of a 10% salt solution should be added to 20 gallons of a 60% salt solution to obtain a 50% salt solution?
Let x be the number of gallons of 10% solution to be added. Then:
.6*20 + .1*x = .5*(20+x)
12 + .1*x = 10 + .5x
.4*x = 2
4*x = 20
x = ?